Darla bicycled at an average rate of 15 miles per hour in a 60 kilometer race. Determine the lenght of time in hours that it took her to complete the race.(hint: 1.6 kilometers= 1 mile)
please help

Answer by stanbon(75887)

Darla bicycled at an average rate of 15 miles per hour in a 60 kilometer race. Determine the lenght of time in hours that it took her to complete the race.(hint: 1.6 kilometers= 1 mile)
-----------------------------------
DATA:
rate = 15 mph = (15*1.6) = 24 km/h
----------------------------------
distance = 60 km
------------------------
time = distance / rate = 60/24 = 2.5 hrs.
================================================
Cheers,
Stan H.